# Unconstrained Open Vocabulary Image Classification: Zero-Shot Transfer from Text to Image via CLIP Inversion

### v1

@@ 3.1 Object Decoder Model
The architecture of the object decoder model is shown in Fig. 3. The model is built around an autoregressive decoder-only language transformer [57] that has been configured to use learned token embeddings, weight tying [47], learned positional encodings, pre-layer normalization [64], GELU activation functions [22], beam search, custom weight initialization, a contractive inner feed-forward dimension, and no biases in the linear and layer normalization layers, including the final token logits linear layer. Dropout is applied after adding the positional encoding and within the transformer layers themselves. The input embedding vector is supplied to the transformer in the form of $P$ prefix token vectors, i.e. $P$ dynamically calculated vectors of dimension $H$ (hidden dimension). The prefix tokens are computed from the input embedding using a learned linear projection that has $PH$ outputs and no biases, as the immediately following positional encoding makes biases redundant. The tokenized target object nouns do not require a start token, as the first output token is predicted at the sequence location of the last prefix token. Appropriate sequence masking is also used to ensure that only sequence locations up to the first output <eos> contribute to the loss. The internal attention mask of the transformer is causal, with the exception that the $P$ prefix tokens can attend to each other without restriction, allowing them to share information freely.
@@ 3.2 Training Dataset for Text-Only Supervision
An exhaustive dictionary of simple and compound object nouns, referred to as the object noun dictionary, was created based on the WordNet dictionary [42], GNU Collaborative International Dictionary of English [5], and the categories used by various image datasets. [^2] A best effort was made to keep the object noun dictionary as accurate and complete as possible given the available dictionary data. A missing word would be indicative of a possible improvement to the object noun dictionary rather than a weakness of the object decoder training scheme. Noun entries were cleaned, collected and merged as appropriate to account for word duplication, plurality, inflection and alternate spellings, and filtered down to only include nouns that correspond to suitable object classifications (e.g. no abstract nouns). This resulted in 42 919 unique case-normalized object nouns corresponding to 96 526 noun variants. Each object noun was then assigned a log-scale frequency value based on its observed frequency in the English language as per the n-gram frequencies in the Google Web Trillion Word Corpus [44, 4]. Object nouns that are categories in a curated, comprehensive collection of image datasets were given appropriate frequency boosts. The log-scale frequencies were scaled and rounded so that very rare nouns saturate at a value of 1, and very frequent nouns have values $\geq$12. Aside from being used to adjust relative sampling frequencies during training, the resulting frequencies allowed the construction of object noun dictionary subsets by filtering out nouns less or equal to a particular frequency threshold (FT), e.g. FT2.333FT0, FT2, FT6, and FT9 have 42 919, 11 897, 5 899, and 2 919 unique object nouns, respectively. FT0 is the complete object noun dictionary.
@@
In order to supply the object decoder with even more diverse synthetic training data to generalize from, captions were generated for each noun in the object noun dictionary using an LLM. This approach leverages the implicit contextual knowledge of LLMs to place object nouns in relevant contexts—incorporating appropriate adjectives, actions, scenes, and secondary objects—in a style resembling real image captions. [^5] CLIP embedding spaces are trained on real image captions. The minimum number of unique generated captions for each noun variant was set proportionally to the corresponding log-scale frequency (up to 100) in order to match the relative word frequencies found in the English language. OpenAI GPT-3.5 Turbo was selected as an ideal balance of cost and sufficient performance, and was used to generate a total of 1.8M captions. Ensuring frequency-appropriate use of alternate spellings, this resulted in a total of 2.9M caption samples (2.0M unique) that were then oversampled to around 30% the size of any multiset they were merged with. For the combined dataset $\mathcal{M}_{1-3}$ this resulted in 51.7M samples with 91.9M targets.
@@
In order to allow more robust and generalizable features of the CLIP embedding space to be learned, data augmentation in the form of noise is used during training. Due to the large modality gap between text and image embeddings [39], noise is in fact a crucial component of the training scheme. Noise allows the object decoder to learn to be invariant to even very large changes in the non-object-related parts of the text embedding vectors, which helps the generalization to image embeddings, as these are very sensitive to secondary objects, backgrounds, and even subtle visual variations. Gu et al. [18] used elementwise Gaussian noise with vector renormalization to train a captioning model on text embeddings only, but did not entirely accurately explain its effect. The CLIP embedding space is deceptively and unimaginably vast [^6] Even if we assumed each element of a 768-dim embedding vector only contains a single bit of information, this would still lead to $10^{231}$ different embeddings, against which even a trillion samples is essentially zero., so the random noise statistically cannot be significantly filling the gaps between samples, or causing a meaningful overlap with image embedding vectors. The main benefit of noise is that the model can learn to focus on which parts (i.e. subspaces) of the embedding vector really encode which individual object concepts, instead of focusing on which region of the embedding space a vector is in as a whole. We observe that the amount of noise required for best training results is actually quite large, not small or minor [18], and its effect is to completely scatter the cone [39] of text embeddings to a thin hyperannulus on the order of 65–75 away, with no single sample remaining significantly closer than that to any original text embedding. By comparison, the very large modality gap between the image and text embedding cones can also be measured, and is around 70–85 for matching image-text pairs (depending on the CLIP model), and only 8–12 more than this for completely uncorrelated image-text pairs. As indicated previously though, statistically there is negligible region overlap between the scattered hyperannulus and the image embeddings cone due to the vast nature of the embedding space. The large amount of noise required can instead be seen as a trade-off between the disruptive ‘blurring’ effect of noise, and the benefit of the object decoder having seen many embedding vectors during training that are as far removed from text embeddings as image embeddings are. In this work, we use an 85% to 15% mix of elementwise Gaussian noise and uniform angle noise. The latter was developed to beneficially cover a greater range of noise dynamics and levels, by rotating input text embedding vectors uniformly in random directions by angles in the uniform 45–75 range.
@@ 3.3 Evaluation Datasets for Zero-Shot Recognition
As NOVIC is trained without providing explicit categories, we wish to evaluate it also on datasets that do not specify categories. No suitable such dataset is known however, as such a dataset would need to have annotations for every possible correct object noun in the English language for every visible entity in every part of each image. To address this, three open vocabulary image datasets were curated as part of this work, and individually annotated by both human and multimodal LLM annotators. The annotations specify whether each classification is correct, close, or incorrect, and for the human annotations, whether it relates to a primary or secondary element of the image. The three new datasets are (i) World: 272 images of which the grand majority are originally sourced (have never been on the internet) from 10 countries by 8 people, with an active focus on covering as wide and varied concepts as possible, including unusual, deceptive and/or indirect representations of objects, (ii) Wiki: 1 000 Wikipedia lead images sampled from a scraped pool of 18K, (iii) Val3K: 3 000 images from the ImageNet-1K validation set, sampled uniformly across the classes (LLM annotations only). The suffixes -H and -L are used to clarify whether human or LLM annotations are in use, e.g. Wiki-H. A total of 16.5K human and 112K LLM (OpenAI GPT-4o) annotations were made across the three datasets. This was at the limit of feasibility in terms of number of samples, person-hours and cost.
@@ 4 Experiments
For experiments, the SigLIP ViT B/16 model [70] is used as the base CLIP model, due to its favorable balance of speed and performance, especially when compared to the original CLIP models [48]. Training is performed for 18 epochs with the AdamW optimizer, cosine learning rate decay, weight decay on all multidimensional weights, and a gradient-accumulated batch size of $B=8192$. The object decoder model is quite small at only 12.2M parameters, and takes less than 3 days to train on a single RTX A6000. For FT2, this drops to 11.4M parameters and only 1.5 days. Overall, this is extremely efficient, considering that up to a billion samples are being seen by a single GPU. As a scale comparison, using our method on the very same GPU to train an object decoder on all 1.28M ImageNet-1K samples for 18 epochs takes a mere 27 minutes, and results in an impressive top-1 accuracy of 83.4%. The object decoder is also very resource efficient at inference time. Requiring at most 5 GB GPU memory, the total pipeline—including CLIP model, autoregression, and beam search—takes on average 26 ms per image for single images, and 7 ms per image when using batching ($B=256$), making it very real-time capable.
@@
Model predictions on the three introduced open vocabulary image datasets are scored by assigning {1, 0.8, 0.5, 0.4, 0} points for {correct primary, correct secondary, close primary, close secondary, incorrect} predictions respectively, and expressing the sum as a percent of the maximum possible score. For verification purposes, the prediction scores achieved by many models were statistically compared between human and LLM annotations on the same dataset and found to be suitably correlated, thus allowing model performance insights to be gained even when manual annotation is just too prohibitive. It is expected that LLM prediction scores are lower than human ones due to some residual LLM mistakes, a generally narrow acceptance of classifications, and inherently less frequent use of the close annotation by the LLM.
@@ 4.1 Zero-Shot Classification Performance
As discussed in Sec. 2, no directly equivalent state-of-the-art model is known that performs real-time prompt-free open vocabulary image classification. The most suitable models for comparison are image tagging models like Tag2Text [24] and the Recognize Anything Model (RAM) [72], as they are designed to be able to provide text labels for images of arbitrary content. We evaluate the largest model variants (14M) on Wiki-H by taking the highest scoring prediction for each image, and compare this to our results in Tab. 3. While the tagging models obtain a high numeric prediction score, a closer look reveals this is due to a severe lack of diversity and specificity (cf. Fig. 4). The produced tags are tendentially coarse and repetitive, with around a third of all images receiving one of only 20 generic top labels, often missing the main point of images. Even when provided with the full object noun dictionary, the total diversity of labels stays below 400, with over 80% of predictions still being made from the small core RAM vocabulary. This is in contrast to our model, which has a much flatter prediction distribution, diversities up to 850 (cf. Tab. 3), and a high prediction score despite generating fine-grained classifications. The tags produced by Tag2Text and RAM also tend to be uninformative—when applied to the images of the Imagewoof [12] validation set for example, both models simply output dog 88% of the time, and never a single dog breed. Our FT0 model, on the other hand, predicts over 90 different dog breeds, with much of the weight covering variations of the actual ground truth classes. When RAM is evaluated on ImageNet-1K by only providing it the 1000 class names to choose from, it only scores 38.3% (25.4% for non-core vocabulary), which is far less than the 75.54% of the ViT-L CLIP model it was trained on, and far less than the 60–70% nominally scored by NOVIC.
@@ 4.2 Ablation Studies
Tab. 4 shows the effect on decoder performance of varying the dataset configuration, training noise, and underlying CLIP model. FT2 was chosen as the ablation baseline, as many ablation runs were required, and FT2 strikes a balance between overall feasibility of training times, and ensuring that a diverse and expressive range of object nouns are seen during training. Overall, as expected, the human annotations exhibit generally higher scores and are slightly more consistent, due to human checking and more prevalent use of the close category. The decoder ImageNet-1K scores are also observed to be well-correlated with general open prediction performance.
@@
Table 4: Dataset, training noise, and CLIP model ablations (mean of 3). IN1K is the ImageNet-1K classification performance of the trained decoders, and the asterisks correspond to the nominal ablation baseline.

### v2

@@ 3.1 Object Decoder Model
The architecture of the object decoder model is shown in Fig. 3. The model is built around an autoregressive decoder-only language transformer [57] that has been configured to use learned token embeddings, weight tying [47], learned positional encodings, pre-layer normalization [64], GELU activation functions [22], beam search, custom weight initialization (cf. App. A), a contractive inner feed-forward dimension, and no biases in the linear and layer normalization layers, including the final token logits linear layer. Dropout is applied after adding the positional encoding and within the transformer layers themselves. The input embedding vector is supplied to the transformer in the form of $P$ prefix token vectors, i.e. $P$ dynamically calculated vectors of dimension $H$ (hidden dimension). The prefix tokens are computed from the input embedding using a learned linear projection that has $PH$ outputs and no biases, as the immediately following positional encoding makes biases redundant. The tokenized target object nouns do not require a start token, as the first output token is predicted at the sequence location of the last prefix token. Appropriate sequence masking is also used to ensure that only sequence locations up to the first output <eos> contribute to the loss. The internal attention mask of the transformer is causal, with the exception that the $P$ prefix tokens can attend to each other without restriction, allowing them to share information freely.
@@ 3.2 Training Dataset for Text-Only Supervision
An exhaustive dictionary of simple and compound object nouns, referred to as the object noun dictionary, was created based on the WordNet dictionary [42], GNU Collaborative International Dictionary of English [5], and the categories used by various image datasets (cf. App. B). [^2] A best effort was made to keep the object noun dictionary as accurate and complete as possible given the available dictionary data. A missing word would be indicative of a possible improvement to the object noun dictionary rather than a weakness of the object decoder training scheme. Noun entries were cleaned, collected and merged as appropriate to account for word duplication, plurality, inflection and alternate spellings, and filtered down to only include nouns that correspond to suitable object classifications (e.g. no abstract nouns). This resulted in 42 919 unique case-normalized object nouns corresponding to 96 526 noun variants. Each object noun was then assigned a log-scale frequency value based on its observed frequency in the English language as per the n-gram frequencies in the Google Web Trillion Word Corpus [44, 4]. Object nouns that are categories in a curated, comprehensive collection of image datasets were given appropriate frequency boosts. The log-scale frequencies were scaled and rounded so that very rare nouns saturate at a value of 1, and very frequent nouns have values $\geq$12. Aside from being used to adjust relative sampling frequencies during training, the resulting frequencies allowed the construction of object noun dictionary subsets by filtering out nouns less or equal to a particular frequency threshold (FT), e.g. FT2.333FT0, FT2, FT6, and FT9 have 42 919, 11 897, 5 899, and 2 919 unique object nouns, respectively. FT0 is the complete object noun dictionary.
@@
In order to supply the object decoder with even more diverse synthetic training data to generalize from, captions were generated for each noun in the object noun dictionary using an LLM (cf. App. C). This approach leverages the implicit contextual knowledge of LLMs to place object nouns in relevant contexts—incorporating appropriate adjectives, actions, scenes, and secondary objects—in a style resembling real image captions. [^5] CLIP embedding spaces are trained on real image captions. The minimum number of unique generated captions for each noun variant was set proportionally to the corresponding log-scale frequency (up to 100) in order to match the relative word frequencies found in the English language. OpenAI GPT-3.5 Turbo was selected as an ideal balance of cost and sufficient performance, and was used to generate a total of 1.8M captions. Ensuring frequency-appropriate use of alternate spellings, this resulted in a total of 2.9M caption samples (2.0M unique) that were then oversampled to around 30% the size of any multiset they were merged with. For the combined dataset $\mathcal{M}_{1-3}$ this resulted in 51.7M samples with 91.9M targets.
@@
In order to allow more robust and generalizable features of the CLIP embedding space to be learned, data augmentation in the form of noise is used during training. Due to the large modality gap between text and image embeddings [39], noise is in fact a crucial component of the training scheme. Noise allows the object decoder to learn to be invariant to even very large changes in the non-object-related parts of the text embedding vectors, which helps the generalization to image embeddings, as these are very sensitive to secondary objects, backgrounds, and even subtle visual variations. Gu et al. [18] used elementwise Gaussian noise with vector renormalization to train a captioning model on text embeddings only, but did not entirely accurately explain its effect. The CLIP embedding space is deceptively and unimaginably vast [^6] Even if we assumed each element of a 768-dim embedding vector only contains a single bit of information, this would still lead to $10^{231}$ different embeddings, against which even a trillion samples is essentially zero., so the random noise statistically cannot be significantly filling the gaps between samples, or causing a meaningful overlap with image embedding vectors. The main benefit of noise is that the model can learn to focus on which parts (i.e. subspaces) of the embedding vector really encode which individual object concepts, instead of focusing on which region of the embedding space a vector is in as a whole. We observe that the amount of noise required for best training results is actually quite large, not small or minor [18], and its effect is to completely scatter the cone [39] of text embeddings to a thin hyperannulus on the order of 66–80 away, with no single sample remaining significantly closer than that to any original text embedding (cf. App. D). By comparison, the very large modality gap between the image and text embedding cones can also be measured, and is around 70–85 for matching image-text pairs (depending on the CLIP model), and only 8–12 more than this for completely uncorrelated image-text pairs. As indicated previously though, statistically there is negligible region overlap between the scattered hyperannulus and the image embeddings cone due to the vast nature of the embedding space. The large amount of noise required can instead be seen as a trade-off between the disruptive ‘blurring’ effect of noise, and the benefit of the object decoder having seen many embedding vectors during training that are as far removed from text embeddings as image embeddings are. In this work, we use an 85% to 15% mix of elementwise Gaussian noise and uniform angle noise. The latter was developed to beneficially cover a greater range of noise dynamics and levels, by rotating input text embedding vectors uniformly in random directions by angles in the uniform 45–75 range.
@@ 3.3 Evaluation Datasets for Zero-Shot Recognition
As NOVIC is trained without providing explicit categories, we wish to evaluate it also on datasets that do not specify categories. No suitable such dataset is known however, as such a dataset would need to have annotations for every possible correct object noun in the English language for every visible entity in every part of each image. To address this, three open vocabulary image datasets were curated as part of this work (cf. App. E), and individually annotated by both human and multimodal LLM annotators (cf. App. F). The annotations specify whether each classification is correct, close, or incorrect, and for the human annotations, whether it relates to a primary or secondary element of the image. The three new datasets are (i) World: 272 images of which the grand majority are originally sourced (have never been on the internet) from 10 countries by 12 people, with an active focus on covering as wide and varied concepts as possible, including unusual, deceptive and/or indirect representations of objects, (ii) Wiki: 1 000 Wikipedia lead images sampled from a scraped pool of 18K, (iii) Val3K: 3 000 images from the ImageNet-1K validation set, sampled uniformly across the classes (LLM annotations only). The suffixes -H and -L are used to clarify whether human or LLM annotations are in use, e.g. Wiki-H. A total of 16.5K human and 112K LLM (OpenAI GPT-4o) annotations were made across the three datasets. This was at the limit of feasibility in terms of number of samples, person-hours and cost.
@@ 4 Experiments
For experiments, the SigLIP ViT B/16 model [70] is used as the base CLIP model, due to its favorable balance of speed and performance, especially when compared to the original CLIP models [48]. Training is performed for 18 epochs with the AdamW optimizer, cosine learning rate decay, weight decay on all multidimensional weights, and a gradient-accumulated batch size of $B=8192$. Hyperparameter details can be found in App. G. The object decoder model is quite small at only 12.2M parameters, and takes less than 3 days to train on a single RTX A6000. For FT2, this drops to 11.4M parameters and only 1.5 days. Overall, this is extremely efficient, considering that up to a billion samples are being seen by a single GPU. As a scale comparison, using our method on the very same GPU to train an object decoder on all 1.28M ImageNet-1K samples for 18 epochs takes a mere 27 minutes, and results in an impressive top-1 accuracy of 83.4%. The object decoder is also very resource efficient at inference time. Requiring at most 5 GB GPU memory, the total pipeline—including CLIP model, autoregression, and beam search—takes on average 26 ms per image for single images, and 7 ms per image when using batching ($B=256$), making it very real-time capable.
@@
Model predictions on the three introduced open vocabulary image datasets are scored by assigning {1, 0.8, 0.5, 0.4, 0} points for {correct primary, correct secondary, close primary, close secondary, incorrect} predictions respectively, and expressing the sum as a percent of the maximum possible score. For verification purposes, the prediction scores achieved by many models were statistically compared between human and LLM annotations on the same dataset and found to be suitably correlated (cf. App. F), thus allowing model performance insights to be gained even when manual annotation is just too prohibitive. It is expected that LLM prediction scores are lower than human ones due to some residual LLM mistakes, a generally narrow acceptance of classifications, and inherently less frequent use of the close annotation by the LLM.
@@ 4.1 Zero-Shot Classification Performance
As discussed in Sec. 2, no directly equivalent state-of-the-art model is known that performs real-time prompt-free open vocabulary image classification. The most suitable models for comparison are image tagging models like Tag2Text [24] and the Recognize Anything Model (RAM) [72], as they are designed to be able to provide text labels for images of arbitrary content. We evaluate the largest model variants (14M) on Wiki-H by taking the highest scoring prediction for each image, and compare this to our results in Tab. 3. While the tagging models obtain a high numeric prediction score, a closer look reveals this is due to a severe lack of diversity and specificity (cf. Figs. 4 and H). The produced tags are tendentially coarse and repetitive, with around a third of all images receiving one of only 20 generic top labels, often missing the main point of images. Even when provided with the full object noun dictionary, the total diversity of labels stays below 400, with over 80% of predictions still being made from the small core RAM vocabulary. This is in contrast to our model, which has a much flatter prediction distribution, diversities up to 850 (cf. Tab. 3), and a high prediction score despite generating fine-grained classifications. The tags produced by Tag2Text and RAM also tend to be uninformative—when applied to the images of the Imagewoof [12] validation set for example, both models simply output dog 88% of the time, and never a single dog breed. Our FT0 model, on the other hand, predicts over 90 different dog breeds, with much of the weight covering variations of the actual ground truth classes. When RAM is evaluated on ImageNet-1K by only providing it the 1000 class names to choose from, it only scores 38.3% (25.4% for non-core vocabulary), which is far less than the 75.54% of the ViT-L CLIP model it was trained on, and far less than the 60–70% nominally scored by NOVIC.
@@ 4.2 Ablation Studies
Tab. 4 shows the effect on decoder performance of varying the dataset configuration, training noise, and underlying CLIP model. FT2 was chosen as the ablation baseline, as many ablation runs were required, and FT2 strikes a balance between overall feasibility of training times, and ensuring that a diverse and expressive range of object nouns are seen during training. Further ablations are presented in App. I. Overall, as expected, the human annotations exhibit generally higher scores and are slightly more consistent, due to human checking and more prevalent use of the close category. The decoder ImageNet-1K scores are also observed to be well-correlated with general open prediction performance.
@@
Table 4: Dataset, training noise, and CLIP model ablations (mean of 3). IN1K is the ImageNet-1K classification performance of the trained decoders, and the asterisks correspond to the nominal ablation baseline. See App. I for more ablation results.